Spectrographs B - E were working. Grating efficiency was very poor in the E spectrograph between about 1300Å to 1500Å.
There was almost no imaging data on this flight because of software problems.
The image intensifier tube on spectrograph C failed on MET day 2. The optics of spectrograph E shifted during launch, so the spectrum from 1150Å to 2000Å was out of focus.
The on-board tape recorder failed, so we were only able to take data at the full rate when the orbiter was in communication with the ground through the ku band. The orbiter was in LOS (Loss of Signal) for roughly 20 percent of the time. After MET day 2, data was taken at a reduced rate during LOS periods and stored in memory until the ku band was restored to us.
There is a 60 hour period during which GLO was making continuous limb observations. The period starts at MET 4/06:04.
On STS-69, the D spectrograph did not work. The optics shifted in the E spectrograph, so the full altitude range is not available in the 1150Å to 2000Å spectral image.
All spectrograph and imager detectors were working on this mission.
There were problems with the tracking which are still being analyzed. There was no night limb tracking, and during the day tracking worked only on a few passes.
A very noisy low-rate communications channel made commanding difficult on this flight.
This page is maintained by: